OpenStudy (anonymous):

Find the vertex quick! Show work please! y=-2^2+4x-3

OpenStudy (anonymous):

what vertex ?

OpenStudy (amistre64):

what is a vertex quick?

OpenStudy (anonymous):

y=-2^2+4x-3 y= 2 to the second power + 4x - 3

OpenStudy (anonymous):

I need to know the vertex

OpenStudy (amistre64):

the vertex is NOT the lowest point on the graph

OpenStudy (amistre64):

it is the point where the graph turns over on itself; kinda like a corner or a bend

OpenStudy (amistre64):

it happens on a quadratic along the axis of symmetry

OpenStudy (anonymous):

using the formula x=-b/2a find the vertex and whether or not it is min or max